# Lightweight posterior construction for gravitational-wave catalogs with the Kolmogorov-Arnold network

### v1

@@ I Introduction
Since the first detection of the gravitational-wave (GW) signal from a binary black hole (BBH) merger Abbott et al. (2016), GWs have opened a new era in our exploration of the Universe. The ground-based GW detectors, LIGO Aasi et al. (2015), Virgo Acernese et al. (2015), and KAGRA Akutsu et al. (2019), have completed three observing runs Abbott et al. (2019a, 2021a, 2024, 2023a), and together with the first part of the fourth run Abac et al. (2025a, b), more than 200 GW events from compact binary coalescences (CBCs) are reported. These observations have driven transformative advances across multiple domains of astrophysics Abbott et al. (2019b, 2021b, 2023b), cosmology Abbott et al. (2017); Soares-Santos et al. (2019); Abbott et al. (2021c, 2023c), and fundamental physics Yagi and Yunes (2013); Abbott et al. (2018, 2021d).
@@
Behind these groundbreaking discoveries lie the advancements in instruments, as well as data analysis techniques. Bayesian inference constitutes the fundamental framework in GW data analysis, providing a principled and systematic approach to parameter estimation and population inference Thrane and Talbot (2019). In Bayesian inference, the parameter estimation is calculating the conditional distribution, the so-called posterior $P(\bm{\theta}|d)$, of the parameters $\bm{\theta}$ given the observed data $d$. By employing Bayes’ theorem, the posterior can be obtained given the prior $\pi(\bm{\theta})$ and the likelihood function $\mathcal{L}(d|\bm{\theta})$, via $P(\bm{\theta}|d)={\mathcal{L}(d|\bm{\theta})\pi(\bm{\theta})}/{\mathcal{Z}(d)}$, where $\mathcal{Z}(d)$ is the normalization factor known as the evidence. Usually, it is unfeasible to directly calculate the exact expression of the posterior. Luckily, with advanced stochastic sampling methods such as the Markov Chain Monte Carlo (MCMC) Christensen and Meyer (1998); Sharma (2017) and the nested sampling Skilling (2004, 2006), we can draw posterior samples from this distribution. Posterior samples constitute key data products in GW catalogs underpinning subsequent scientific investigation in GW astronomy.
@@
As GW detections become routine, the focus of GW astrophysics is expanding from analyzing individual events to conducting systematic studies on the entire population of sources. This shift necessitates a compilation of comprehensive and accessible GW catalogs. Growing catalogs call for powerful statistical techniques, like the hierarchical Bayesian analysis Thrane and Talbot (2019); Wang et al. (2024), to unravel the collective properties of compact binaries. Distributions of compact binary source parameters, such as masses Abbott et al. (2023b); Golomb and Talbot (2022); Dong et al. (2024) and spins Abbott et al. (2023b); Zhu et al. (2018); Tong et al. (2022) provide critical constraints on models of stellar evolution and compact object formation. Furthermore, combining information from multiple events allows for stringent tests of Einstein’s theory of general relativity Abbott et al. (2021d) and precise constraints on the properties of dense matter, such as the equation of state of neutron stars Landry and Essick (2019); Wysocki et al. (2020); Shao and Yagi (2022); Wang et al. (2024). Therefore, efficiently constructing and utilizing GW catalogs is crucial to advance our understanding of the Universe, especially in the era of next-generation GW detectors when many more events are detected.
@@
Next-generation ground-based GW detectors, such as the Einstein Telescope (ET) Punturo et al. (2010); Branchesi et al. (2023); Abac et al. (2025c) and Cosmic Explorer (CE) Reitze et al. (2019); Evans et al. (2021), will achieve unprecedented sensitivities, unlocking new scientific opportunities Sathyaprakash et al. (2019); Kalogera et al. (2021). They will boost the detection rates for binary neutron-star and black-hole mergers, reaching $10^{5}$$10^{6}$ events per year Samajdar et al. (2021); Himemoto et al. (2021); Hu and Veitch (2023). Considering that standard offline Bayesian inference for a single event already demands high computational resources, the high event rates expected in the next-generation GW detectors will introduce new computational challenge for data analysis. As predicted, performing standard Bayesian inference on a one-month catalog requires $10^{13}$$10^{15}$ CPU hours, imposing a substantial burden on computing infrastructure, electricity cost, and environmental impact Hu and Veitch (2024). Efficient, accelerated Bayesian methods are therefore urgently needed. Besides, from the perspective of data storage and transmission, posterior-sample-based GW catalogs with millions of events could easily reach a terabyte-scale volume, placing significant burdens on users for both transmission and storage.
@@
The development of machine learning techniques offers a possible solution to the data challenges posed by next-generation GW detectors. Machine learning has already demonstrated its great potential in GW detection George and Huerta (2018); Huerta et al. (2018); Xia et al. (2021); Ma et al. (2023, 2025); Wang and Zeng (2025), parameter estimation Gabbard et al. (2022); Green et al. (2020); D’Emilio et al. (2021); Santoliquido et al. (2025); Green and Gair (2021); Dax et al. (2021, 2025), and subsequent astrophysical researches Leyde et al. (2024); McEwen et al. (2023); Talbot and Thrane (2022); Mould et al. (2022); Ventagli and Saltas (2025); Tang and Fan (2024); Hu (2025). Specifically, neural posterior estimation has been widely adopted across multiple GW Bayesian inference tasks Dax et al. (2021, 2025); Leyde et al. (2024); McEwen et al. (2023), accelerating the analysis by orders of magnitude. In particular, neural posterior estimation with normalizing flows Green and Gair (2021); Dax et al. (2021, 2025); Leyde et al. (2024); Langendorff et al. (2023); Kobyzev et al. (2020) achieves real-time Bayesian inference for binary black hole mergers and binary neutron star mergers, opening new avenues for multi-messenger astronomy. In the era of next-generation detectors, neural posterior estimation has also been explored to address data challenges such as overlapping signals Langendorff et al. (2023); Hu (2025) and long-duration signal analysis Hu et al. (2025). For neural posterior estimation, it needs a neural-network-based density estimator to represent distributions. These neural density estimators are usually built on generative-model architectures, such as variational autoencoders Kingma and Welling (2013); Gabbard et al. (2022), normalizing flows, and flow matching frameworks Chen et al. (2018); Lipman et al. (2023); Liang et al. (2024). For instance, normalizing flows learn a sequence of invertible transformations that map a simple base distribution to the complex target distribution, enabling both fast sampling and exact density evaluation. By efficiently capturing high-dimensional, multimodal distributions, neural density estimators show great promise for alleviating the computational bottlenecks in GW Bayesian inference.
@@
In this work, we propose a Kolmogorov-Arnold network (KAN) Liu et al. (2025a, 2024) based neural density estimator and explore its potential for constructing high-fidelity, low-storage surrogate GW catalogs. KAN is a machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem Arnold (2009); Braun and Griebel (2009). Rather than using fixed activations in conventional multi-layer perceptrons (MLPs), KAN employs edge-wise learnable activation functions. This modification enables KAN to achieve superior interpretability, higher accuracy, and greater parameter efficiency compared to conventional architectures on small-scale artificial intelligence (AI) for scientific tasks. Furthermore, this design also allows the network to converge toward interpretable analytic expressions, a process referred to as the symbolification of the neural network. Currently, KANs are used in fields such as fluid dynamics Toscano et al. (2025); Wang et al. (2025), nuclear physics Liu et al. (2025b), complex dynamical systems Panahi et al. (2025), molecular property prediction Li et al. (2025), and cosmology Cui et al. (2025), and KAN-based machine learning architectures are also rapidly advancing Yang and Wang (2025); Fang et al. (2025); Huang et al. (2025). The rapid development of KAN offers a promising approach to enhance the interpretability of neural networks in AI for science tasks. In our case, to support downstream studies based on GW posteriors, one can download compact, encoded files from our surrogate GW catalogs and rapidly resample the posterior distributions locally. This strategy enables efficient catalog construction while dramatically reducing the demands on data transmission and storage.
@@ II KAN-based neural density estimator
In this section, we introduce the method to compress the posterior samples through a KAN-based neural density estimator. First, we introduce KAN Liu et al. (2025a, 2024) in Sec. II.1. We explain how KAN improves upon conventional neural networks and enables symbolification, after which the entire model is represented merely by the analytic expressions of its output nodes. In Sec. II.2, we present the Masked Autoencoder for Distribution Estimation (MADE) Germain et al. (2015), one of the core building blocks commonly employed in normalizing flows Papamakarios et al. (2017). MADE serves as our base density estimator architecture, which we then enhance using the KAN framework. Section II.3 describes how to perform rapid resampling using the neural network weights and/or the analytic expressions. We outline the workflow for lightweight GW catalog construction in Fig. 1.
@@ II.1 Kolmogorov-Arnold Network
KAN is a novel machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem Arnold (2009); Braun and Griebel (2009). Unlike conventional MLPs that rely on fixed activation functions at each neuron, KAN introduces learnable activation functions on the edges connecting neurons between layers. These edge-wise functions are parameterized as a linear combination of B-spline curves Gordon and Riesenfeld (1974), with coefficients being learnable parameters that are optimized during training. This design allows KAN to capture complex nonlinear relationships between inputs and outputs more flexibly, leading to enhanced accuracy and greater parameter efficiency, particularly in small-scale scientific AI applications.
A key feature of KAN is its capability for network symbolification, a two-stage process involving function matching and synthesis. First, the trained B-spline activation functions on each edge are matched to a predefined library of elementary mathematical functions (e.g., polynomial, trigonometric, and exponential functions) to derive their analytic representations. Subsequently, these individual analytic expressions are composed according to the network’s topology, yielding closed-form analytic expressions that describe the network’s outputs as explicit functions of its inputs Liu et al. (2025a). This process significantly enhances the interpretability of the neural network, making its internal mechanism more transparent and analyzable. Leveraging KAN’s efficient function representation and unique symbolification capabilities, we propose a KAN-based neural density estimator. Our objective is to harness these advantages to achieve substantial storage compression of GW posterior samples, thereby offering a new approach for constructing lightweight and user-friendly GW catalogs.
@@ II.2 Masked Autoencoder for Distribution Estimation
To model the probability distribution of the posterior samples, we employ the MADE framework. MADE facilitates density estimation by decomposing the joint probability distribution autoregressively Germain et al. (2015). For a $D$-dimensional parameter vector $\bm{\theta}$, this decomposition is expressed as
@@
where $\bm{\theta}_{<i}$ denotes the set of parameters $\{\theta_{1},\theta_{2},\cdots,\theta_{i-1}\}$. MADE approximates each conditional probability $P_{i}$ with a Gaussian mixture model (GMM) McEwen et al. (2023); Germain et al. (2015); Papamakarios and Murray (2016), such that
@@
A neural network, denoted by its weights $w$, is trained to output the parameters of these GMMs. To ensure positivity of $a_{ij}$ and $\sigma_{ij}$, the network predicts $\ln(a_{ij})$ and $2\ln(\sigma_{ij})$ alongside $\mu_{ij}$. The complete network-approximated distribution is denoted as $P_{w}(\bm{\theta})$. The crucial autoregressive property is enforced by applying binary masks to the network’s connections, ensuring that each conditional probability $P_{i}$ depends only on $\bm{\theta}_{<i}$. Alternative approaches for enforcing autoregressive constraints include masked convolutions Van Den Oord et al. (2016) and causal convolutions van den Oord et al. (2016). If the target distribution is conditional $P(\bm{\theta}|\bm{\phi})$, the conditioning variables $\bm{\phi}$ need to be connected to the network through unmasked connections.
@@ II.3 Resampling Using Neural Network Weights and Analytic Expressions
For the resampling with neural network weights, the sample generation process is accomplished iteratively according to Eq. (1) where $\bm{\theta}$ is set to $\bm{0}$ in the beginning. During the $i$-th iteration, $\bm{\theta}_{<i}$ is already generated and $\bm{\theta}_{\geq i}$ is still $\bm{0}$. Then $\{\bm{\theta}_{<i},\bm{0}\}$ is fed into the network to obtain $a_{i,1\leq j\leq G}$, $\mu_{i,1\leq j\leq G}$, and $\sigma_{i,1\leq j\leq G}$. Next, $\theta_{i}$ is sampled from the conditional distribution $P_{i}(\theta_{i}|\bm{\theta}_{<i})$. However, in each sequential iteration, all network outputs are computed, introducing computational inefficiency as many node computations—those not related to $P_{i}$—are unnecessary. If we assume that the network does not contain hidden layers (as we do), then the computational complexity of resampling is $\mathcal{O}\,\big{(}GD^{3}\big{)}$. This is because that the number of nodes in the input layer is $D$, the number of nodes in the output layer is proportional to $GD$, and the number of iterations equals $D$. Probability density computation using a neural network requires one forward computation to get all the GMM parameters and the complexity is $\mathcal{O}\,\big{(}GD^{2}\big{)}$. Then $P_{w}(\bm{\theta})$ is computed according to Eq. (1) and Eq. (2).

### v2

@@ I Introduction
Since the first detection of the gravitational-wave (GW) signal from a binary black hole (BBH) merger LIGOScientific:2016aoc, GWs have opened a new era in our exploration of the Universe. The ground-based GW detectors, LIGO LIGOScientific:2014pky, Virgo VIRGO:2014yos, and KAGRA KAGRA:2018plz, have completed three observing runs LIGOScientific:2018mvr; LIGOScientific:2020ibl; LIGOScientific:2021usb; KAGRA:2021vkt, and together with the first part of the fourth run LIGOScientific:2025hdt; LIGOScientific:2025slb, more than 200 GW events from compact binary coalescences (CBCs) are reported. These observations have driven transformative advances across multiple domains of astrophysics LIGOScientific:2018jsj; LIGOScientific:2020kqk; KAGRA:2021duu, cosmology ligo2017gravitational; DES:2019ccw; LIGOScientific:2019zcs; LIGOScientific:2021aug, and fundamental physics Yagi:2013awa; LIGOScientific:2018cki; LIGOScientific:2021sio.
@@
Behind these groundbreaking discoveries lie the advancements in instruments, as well as data analysis techniques. Bayesian inference constitutes the fundamental framework in GW data analysis, providing a principled and systematic approach to parameter estimation and population inference thrane2019introduction. In Bayesian inference, the parameter estimation is calculating the conditional distribution, the so-called posterior $P(\bm{\theta}|d)$, of the parameters $\bm{\theta}$ given the observed data $d$. By employing Bayes’ theorem, the posterior can be obtained given the prior $\pi(\bm{\theta})$ and the likelihood function $\mathcal{L}(d|\bm{\theta})$, via $P(\bm{\theta}|d)={\mathcal{L}(d|\bm{\theta})\pi(\bm{\theta})}/{\mathcal{Z}(d)}$, where $\mathcal{Z}(d)$ is the normalization factor known as the evidence. Usually, it is unfeasible to directly calculate the exact expression of the posterior. Luckily, with advanced stochastic sampling methods such as Markov Chain Monte Carlo (MCMC) Christensen:1998gf; Sharma:2017wfu and nested sampling Skilling:2004pqw; Skilling:2006gxv, we can draw posterior samples from this distribution. Posterior samples constitute key data products in GW catalogs underpinning subsequent scientific investigation in GW astronomy.
@@
As GW detections become routine, the focus of GW astrophysics is expanding from analyzing individual events to conducting systematic studies on the entire population of sources. This shift necessitates a compilation of comprehensive and accessible GW catalogs. Growing catalogs call for powerful statistical techniques, like hierarchical Bayesian analyses Mandel:2018mve; thrane2019introduction; Vitale:2020aaz; Wang:2024xon; Gerosa:2024isl, to unravel the collective properties of compact binaries. Distributions of compact binary source parameters, such as masses KAGRA:2021duu; Golomb:2021tll; Dong:2024zzl and spins KAGRA:2021duu; Zhu:2017znf; Tong:2022iws provide critical constraints on models of stellar evolution and compact object formation. Furthermore, combining information from multiple events allows for stringent tests of Einstein’s theory of general relativity LIGOScientific:2021sio and precise constraints on the properties of dense matter, such as the equation of state of neutron stars Landry:2018prl; Wysocki:2020myz; Shao:2022koz; Wang:2024xon. Therefore, efficiently constructing and utilizing GW catalogs is crucial to advance our understanding of the Universe, especially in the era of next-generation GW detectors when many more events are detected.
@@
Next-generation ground-based GW detectors, such as the Einstein Telescope (ET) Punturo:2010zz; Branchesi:2023mws; Abac:2025saz and Cosmic Explorer (CE) Reitze:2019iox; Evans:2021gyd, will achieve unprecedented sensitivities, unlocking new scientific opportunities Sathyaprakash:2019yqt; Kalogera:2021bya. They will boost the detection rates for binary neutron-star and black-hole mergers, reaching $10^{5}$$10^{6}$ events per year Himemoto:2021ukb; Samajdar:2021egv; Hu:2022bji; Borhanian:2022czq; Pieroni:2022bbh; Ronchini:2022gwk; Iacovelli:2022bbs; Begnoni:2025oyd. Considering that standard offline Bayesian inference for a single event already demands high computational resources, the high event rates expected in the next-generation GW detectors will introduce new computational challenge for data analysis. As predicted, performing standard Bayesian inference on a one-month catalog requires $10^{13}$$10^{15}$ CPU hours, imposing a substantial burden on computing infrastructure, electricity cost, and environmental impact Hu:2024mvn. Efficient, accelerated Bayesian methods are therefore urgently needed. Besides, from the perspective of data storage and transmission, posterior-sample-based GW catalogs with millions of events could easily reach a terabyte-scale volume, placing significant burdens on users for both transmission and storage.
@@
The development of machine learning techniques offers a possible solution to the data challenges posed by next-generation GW detectors. Machine learning has already demonstrated its great potential in GW detection George:2017pmj; Huerta:2017kez; Xia:2020vem; Ma:2023ctz; Ma:2024log; Wang:2025xvj, parameter estimation Gabbard:2019rde; Green:2020hst; DEmilio:2021laf; Santoliquido:2025lot; Green:2020dnx; Dax:2021tsq; Dax:2024mcn, and subsequent astrophysical researches Leyde:2023iof; mcewen2021machine; Talbot:2020oeu; Mould:2022ccw; Ventagli:2024xsh; Tang:2024orq; Hu:2025vlp. Specifically, neural posterior estimation has been widely adopted across multiple GW Bayesian inference tasks Dax:2021tsq; Dax:2024mcn; Leyde:2023iof; mcewen2021machine, accelerating the analysis by orders of magnitude. In particular, neural posterior estimation with normalizing flows Green:2020dnx; Dax:2021tsq; Dax:2024mcn; Leyde:2023iof; Langendorff:2022fzq; kobyzev2020normalizing achieves real-time Bayesian inference for binary black hole mergers and binary neutron star mergers, opening new avenues for multi-messenger astronomy. In the era of next-generation detectors, neural posterior estimation has also been explored to address data challenges such as overlapping signals Langendorff:2022fzq; Hu:2025vlp and long-duration signal analysis Hu:2024lrj. For neural posterior estimation, it needs a neural-network-based density estimator to represent distributions. These neural density estimators are usually built on generative-model architectures, such as variational autoencoders Kingma:2013hel; Gabbard:2019rde, normalizing flows, and flow matching frameworks Chen:2018wjc; lipman2023flow; Liang:2024new. For instance, normalizing flows learn a sequence of invertible transformations that map a simple base distribution to the complex target distribution, enabling both fast sampling and exact density evaluation. By efficiently capturing high-dimensional, multimodal distributions, neural density estimators show great promise for alleviating the computational bottlenecks in GW Bayesian inference.
@@
In this work, we propose a Kolmogorov-Arnold network (KAN) liu2025kan; liu2024kan2 based neural density estimator and explore its potential for constructing high-fidelity, low-storage surrogate GW catalogs. KAN is a machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem arnold2009representation; braun2009constructive. Rather than using fixed activations in conventional multi-layer perceptrons (MLPs), KAN employs edge-wise learnable activation functions. This modification enables KAN to achieve superior interpretability, higher accuracy, and greater parameter efficiency compared to conventional architectures on small-scale artificial intelligence (AI) for scientific tasks. Furthermore, this design also allows the network to converge toward interpretable analytic expressions, a process referred to as the symbolification of the neural network. Currently, KANs are used in fields such as fluid dynamics Toscano2025; Wang_2025, nuclear physics PhysRevC.111.024316, complex dynamical systems PhysRevResearch.7.023037, molecular property prediction li2025KAGNN, and cosmology Cui:2025rri; Jones:2025obf, and KAN-based machine learning architectures are also rapidly advancing yang2025kolmogorovarnold; fang2025kaa; huang2025timekan. The rapid development of KAN offers a promising approach to enhance the interpretability of neural networks in AI for science tasks. In our case, to support downstream studies based on GW posteriors, one can download compact, encoded files from our surrogate GW catalogs and rapidly resample the posterior distributions locally. This strategy enables efficient catalog construction while dramatically reducing the demands on data transmission and storage.
@@ II KAN-based neural density estimator
In this section, we introduce the method to compress the posterior samples through a KAN-based neural density estimator. First, we introduce KAN liu2025kan; liu2024kan2 in Sec. II.1. We explain how KAN improves upon conventional neural networks and enables symbolification, after which the entire model is represented merely by the analytic expressions of its output nodes. In Sec. II.2, we present the Masked Autoencoder for Distribution Estimation (MADE) germain2015made, one of the core building blocks commonly employed in normalizing flows Papamakarios:2017tec. MADE serves as our base density estimator architecture, which we then enhance using the KAN framework. Section II.3 describes how to perform rapid resampling using the neural network weights and/or the analytic expressions. We outline the workflow for lightweight GW catalog construction in Fig. 1.
@@ II.1 Kolmogorov-Arnold Network
KAN is a novel machine-learning architecture inspired by the Kolmogorov-Arnold representation theorem arnold2009representation; braun2009constructive. Unlike conventional MLPs that rely on fixed activation functions at each neuron, KAN introduces learnable activation functions on the edges connecting neurons between layers. These edge-wise functions are parameterized as a linear combination of B-spline curves gordon1974b, with coefficients being learnable parameters that are optimized during training. This design allows KAN to capture complex nonlinear relationships between inputs and outputs more flexibly, leading to enhanced accuracy and greater parameter efficiency, particularly in small-scale scientific AI applications.
A key feature of KAN is its capability for network symbolification, a two-stage process involving function matching and synthesis. First, the trained B-spline activation functions on each edge are matched to a predefined library of elementary mathematical functions (e.g., polynomial, trigonometric, and exponential functions) to derive their analytic representations. Subsequently, these individual analytic expressions are composed according to the network’s topology, yielding closed-form analytic expressions that describe the network’s outputs as explicit functions of its inputs liu2025kan. This process significantly enhances the interpretability of the neural network, making its internal mechanism more transparent and analyzable. Leveraging KAN’s efficient function representation and unique symbolification capabilities, we propose a KAN-based neural density estimator. Our objective is to harness these advantages to achieve substantial storage compression of GW posterior samples, thereby offering a new approach for constructing lightweight and user-friendly GW catalogs.
@@ II.2 Masked Autoencoder for Distribution Estimation
To model the probability distribution of the posterior samples, we employ the MADE framework. MADE facilitates density estimation by decomposing the joint probability distribution autoregressively germain2015made. For a $D$-dimensional parameter vector $\bm{\theta}$, this decomposition is expressed as
@@
where $\bm{\theta}_{<i}$ denotes the set of parameters $\{\theta_{1},\theta_{2},\cdots,\theta_{i-1}\}$. MADE approximates each conditional probability $P_{i}$ with a Gaussian mixture model (GMM) mcewen2021machine; germain2015made; papa2016fast, such that
@@
A neural network, denoted by its weights $w$, is trained to output the parameters of these GMMs. To ensure positivity of $a_{ij}$ and $\sigma_{ij}$, the network predicts $\ln(a_{ij})$ and $2\ln(\sigma_{ij})$ alongside $\mu_{ij}$. The complete network-approximated distribution is denoted as $P_{w}(\bm{\theta})$. The crucial autoregressive property is enforced by applying binary masks to the network’s connections, ensuring that each conditional probability $P_{i}$ depends only on $\bm{\theta}_{<i}$. Alternative approaches for enforcing autoregressive constraints include masked convolutions van2016pixel and causal convolutions oord2016wavenet. If the target distribution is conditional $P(\bm{\theta}|\bm{\phi})$, the conditioning variables $\bm{\phi}$ need to be connected to the network through unmasked connections.
@@ II.3 Resampling Using Neural Network Weights and Analytic Expressions
For the resampling with neural network weights, the sample generation process is accomplished iteratively according to Eq. (1) where $\bm{\theta}$ is set to $\bm{0}$ in the beginning. During the $i$-th iteration, $\bm{\theta}_{<i}$ is already generated and $\bm{\theta}_{\geq i}$ is still $\bm{0}$. Then $\{\bm{\theta}_{<i},\bm{0}\}$ is fed into the network to obtain $a_{i,1\leq j\leq G}$, $\mu_{i,1\leq j\leq G}$, and $\sigma_{i,1\leq j\leq G}$. Next, $\theta_{i}$ is sampled from the conditional distribution $P_{i}(\theta_{i}|\bm{\theta}_{<i})$. However, in each sequential iteration, all network outputs are computed, introducing computational inefficiency as many node computations—those not related to $P_{i}$—are unnecessary. If we assume that the network does not contain hidden layers (as we do), then the computational complexity of resampling is $\mathcal{O}\,\big(GD^{3}\big)$. This is because if the number of nodes in the input layer is $D$, the number of nodes in the output layer is proportional to $GD$, and the number of iterations equals $D$. Probability density computation using a neural network requires one forward computation to get all the GMM parameters and the complexity is $\mathcal{O}\,\big(GD^{2}\big)$. Then $P_{w}(\bm{\theta})$ is computed according to Eq. (1) and Eq. (2).